# Optimizing Layout of Recursive Datatypes with Marmoset

### v2

@@ Abstract
While programmers know that the low-level memory representation of data structures can have significant effects on performance, compiler support to optimize the layout of those structures is an under-explored field. Prior work has optimized the layout of individual, non-recursive structures without considering how collections of those objects in linked or recursive data structures are laid out.
@@ 2.2 Running Example
A non-empty blog value stores a content field (a string representing the body of the blog post), a list of hash tags summarizing keywords of the blog post, and a pointer to the rest of the list. [^2] In practice, you may want to reuse a standard list type, e.g.: type BlogList = [Blog] data Blog = Blog Content HashTags but a typical compiler (including Gibbon) would specialize the parametric list type with the Blog type to arrive at an equivalent of the definition shown in the main text above.. The datatype has one point of recursion and several variable-length fields in the definition. To extend on this, Section 5.3.4 contains an example of tree-shaped data (two points of recursion, in particular) with a fixed-length field. The most general case of multiple points of recursion and variable-length fields is also handled by Marmoset.
@@
Marmoset, described in the following section, automates the process of finding the weaknesses of the pointer-based approach and improving the layout and code accordingly. For instance, performance in our example can be improved by swapping the ordering between the Content and HashTags fields. Given this reordering, the hash tags are available directly at the start of the value, which lines up with the algorithm better, as the algorithm always accesses this field first. Additionally, our program’s True case (the keyword gets a hit within the hash tags) is more efficient because after traversing the content to highlight the keyword it stops at the next blog entry ready for the algorithm to make the recursive call. This improved data layout results in the more-streamlined access patterns shown in Figures 3(b) and 3(d).
@@ 3.1 Marmoset’s Language
Marmoset operates on the language $\lambda_{M}$ given in Figure 4. $\lambda_{M}$ is a first-order, monomorphic, call-by-value functional language with algebraic datatypes and pattern matching. Programs consist of a series of datatype definitions, function definitions, and a main expression. $\lambda_{M}$’s expressions use A-normal form [12]. The notation $\overline{x}$ denotes a vector $[x_{1},\ldots,x_{n}]$ and $\overline{x_{\mathit{i}}}$ the item at position $\mathit{i}$. $\lambda_{M}$ is an intermediate representation (IR) used towards the front end in the Gibbon compiler. The monomorphizer and specializer lower a program written in a polymorphic, higher-order subset of Haskell [^3] With strict evaluation semantics using -XStrict. to $\lambda_{M}$, and then location inference is used to convert it to the location calculus (LoCal) code next [23]. It is easier to update the layout of types in $\lambda_{M}$ compared to LoCal, as in $\lambda_{M}$ the layout is implicitly determined by the ordering of fields, whereas the later LoCal IR makes the layout explicit using locations and regions (essentially, buffers and pointer arithmetic).
@@ 3.2 Control-Flow Analysis
We construct a control-flow graph with sub-expressions, and let-bound RHS’s (right hand sides) of $\lambda_{M}$ as the nodes. Algorithm 1 shows the psuedocode for generating the control-flow graph. Because the syntax is flattened into A-normal form, there is no need to traverse within the RHS of a let expression. Edges between the nodes represent paths between expressions. The edges consist of weights (Line 11) that represent the likelihood of a particular path being taken. An edge between two nodes indicates the order of the evaluation of the program. A node corresponding to a let-binding (Line 9) contains the bound expression and has one outgoing edge to a node corresponding to the body expression. A case expression (Line 14) splits the control flow $n$-ways, where $n$ is the number of pattern matches. Outgoing edges of a node for a case expression have weights associated with them that correspond to the likelihood of taking a particular branch in the program. If available, dynamic profiling information can provide these weights. For the purposes of this paper, however, we find it is sufficient to use trivial placeholder weights: i.e. uniform weight over all branches. Control flow terminates on a variable reference $\lambda_{M}$ expression.
Figure 5(a) shows the control-flow graph for the running example from Figure 1. Each node corresponds to a sub-expression of the function emphKeyword. The first case expression splits the control flow into two branches, corresponding to whether the input list of blogs is empty or not. The branch corresponding to the empty input list is assigned a probability $\alpha$, and thus, the other branch is assigned a probability $1-\alpha$. The next node corresponds to the pattern match Blog content hashTags blogs'. Another two-way branch follows, corresponding to whether keyword occurs in the content of this blog or not. We assign the probabilities $\sigma$ and $1-\sigma$ to these branches respectively. Note that as a result, the corresponding edges in the CFG have weights $(1-\alpha)*\sigma$ and $(1-\alpha)*(1-\sigma)$, as the likelihood of reaching that condition is $(1-\alpha)$. Each of these branches terminate by creating a new blog entry with its content potentially updated. In the current model, $\alpha$ and $\sigma$ are 0.5: they are uniformly distributed.
@@
Figure 5: Control-flow and corresponding field access graphs generated for the running example
@@ 3.4 Field Attributes For Code Motion
When trying to find the best layout, we may treat the code as immutable, but allowing ourselves to move the code around (i.e. change the order of accesses to the fields) unlocks more possibilities for the resulting layout. Not all code motions are valid due to existing data dependencies in the traversal. For instance, in a sequence of two let binders, the second one may reference the binding introduced in the first one: in this case, the two binders cannot be reordered.
@@
A single field can have multiple attributes. For instance, a function (traverse) doing a pre-order traversal on a Tree datatype makes recursive calls on both the left and right children. The left and right children are recursive and self-recursive. Therefore, when looking at the scope of traverse, the left and right children have the attributes recursive, self-recursive and inlineable.
@@
Example. Consider the example of a list traversal shown in Figure 6(a). Here the function foo does some work on the Int field (it raises the Int to the power of 100) and then recurs on the tail of the list. Note that Marmoset is compiling with dense representations, i.e., List type is going to be packed in memory with a representation that stores the Cons tag (one byte) followed by the Int (8 bytes) followed by the next Cons tag and so on. Hence, the Cons tag and the Int field are interleaved together in memory. The function foo becomes a stream processor that consumes one stream in memory and produces a dense output buffer of the same type.
On the other hand, the other way to lay out the list is to change its representation to the following.
@@ 3.5 Field Access Pattern Analysis
18:             mutate edges'= addEdge(edges', ((lastAccessedVar, var), weight), CfgTag)
@@
20:                 mutate edges'= addEdge(edges', ((lastAccessedVar, var), weight), DfgTag)
@@ 3.6.4 Greedy layout ordering
Since there is a tradeoff in finding optimal layout orders and compile times when using a solver based optimization. We implemented a greedy algorithm that traverses the access graph for a data constructor in a greedy fashion. It starts from the root node of the access graph which is the first field in the layout of the data constructor and greedily (based on the edge weight) visits the next field. We fix the edge order for a control-flow edge as the original order and do not look at field attributes. However, after the greedy algorithm picks a layout we match the let expressions to the layout order to make sure the code matches the layout order. A greedy algorithm is potentially sub-optimal when it comes to finding the best performing layout; however, the compile time is fast.
@@ 3.8 Finding a layout for functions with conflicting access patterns
We have conflicting access patterns for D if there are two functions F1 and F2 that access the fields of D in two opposite orders. Assume F1 accesses A first and then B, while F2 accesses B first and then A. After combining edges across the two functions, we will have two edges that are in opposition to each other. Since we give a uniform weight to each function, the edges will also have a uniform weight. In this case, placing A before B or B before A are equally acceptable in our model, and Marmoset defers to the solver to get one of the two possible layouts.
With the two equally good layouts, Marmoset’s default solver-based mode chooses the layout favoring the function it picked first. For instance, if F1 is defined earlier in the program, the order favoring F1 will be picked. In the greedy-heuristic mode, since both A and B are root nodes, Marmoset will pick the first root node in the list of root nodes and start greedily visiting child nodes from that particular root node, thus, fixing a layout order. If we determined that F2 has a greater weight than F1, then the layout preferred by F2 would be chosen.
@@ 5 Evaluation
We evaluate Marmoset on three applications. First is a pair of microbenchmarks: a list length function (Section 5.2.1) and a logical expression evaluator (Section 5.2.2). Second is a small library of operations with binary trees (Section 5.3). Third is a blog management software based on the BlogList example from the Sections 23 (Section 5.4). Besides the run times, we take a closer look at how Marmoset affects cache behavior (Section 5.5) and compile times (Section 10).
We detail the impact of various datatype layouts on the performance. As the baseline, we use Gibbon, the most closely related prior work. We also compare Marmoset with MLton (Section 5.4.5). For each benchmark, we run 99 iterations and report the run-time mean, median and the 95% confidence interval.
@@ 5.2 Micro Benchmarks
We use two microbenchmarks to explore specific performance penalties imposed by a sub-optimal data layout.
5.2.1 Computing the length of a linked-list
This benchmark demonstrates the cost of de-referencing memory addresses which are not present in the cache. It uses the linked list datatype:
@@ 5.2.1 Computing the length of a linked-list
In our experiment, the linked list is made of 3M elements and each element contains an instance of the Pandoc Inline datatype that occupies roughly 5KB. As seen in table 1(a) [^7] The performance of List' layout compiled through Gibbon differs from the greedy and solver version as Marmoset does code motion to reorder let expressions which results in different code., the performance of the list constructed using the original List is $\sim$60$\times$ worse than the performance with the Marmoset-optimized, flipped layout. Not only does List have poor data locality and cache behavior, but it also has to execute more instructions to de-reference the pointer. Both Mgreedy and Msolver choose the flipped layout List'.
Table 1: Mean, Median and 95% confidence interval (ub, lb) for different layouts when compiled with Gibbon. The last two columns show the runtime for the layout chosen by Mgreedy and Msolver. The numbers in Blue correspond to the lowest running time and the numbers in Red correspond to the highest running time. Legend: l - left subtree, r - right subtree of the mono tree.
(a) List length run times in milliseconds (ms)
Gibbon Marmoset
@@
Mean 60.731 1.052 0.981 0.994
Median 60.730 1.017 0.978 0.979
ub 60.737 1.078 0.990 1.022
lb 60.725 1.025 0.973 0.964
(b) Logical expression evaluator run times in seconds (s)
Gibbon Marmoset
@@
Mean 4.372 658 620 370 3714.3726586203703714.372\,658\,620\,370\,3714.372 658 620 370 371 6.669 544 222 222 2246.6695442222222246.669\,544\,222\,222\,2246.669 544 222 222 224 3.600 981 037 037 0373.6009810370370373.600\,981\,037\,037\,0373.600 981 037 037 037 3.594 482 435 185 1853.5944824351851853.594\,482\,435\,185\,1853.594 482 435 185 185
Median 4.313 4384.3134384.313\,4384.313 438 6.625 1076.6251076.625\,1076.625 107 3.534 9583.5349583.534\,9583.534 958 3.543 093 500 000 000 33.54309350000000033.543\,093\,500\,000\,000\,33.543 093 500 000 000 3
ub 4.403 732 645 795 3534.4037326457953534.403\,732\,645\,795\,3534.403 732 645 795 353 6.693 175 524 408 939 56.69317552440893956.693\,175\,524\,408\,939\,56.693 175 524 408 939 5 3.648 194 996 569 0513.6481949965690513.648\,194\,996\,569\,0513.648 194 996 569 051 3.635 610 184 890 1183.6356101848901183.635\,610\,184\,890\,1183.635 610 184 890 118
lb 4.341 584 594 945 3884.3415845949453884.341\,584\,594\,945\,3884.341 584 594 945 388 6.645 912 920 035 5086.6459129200355086.645\,912\,920\,035\,5086.645 912 920 035 508 3.553 767 077 505 0233.5537670775050233.553\,767\,077\,505\,0233.553 767 077 505 023 3.553 354 685 480 2523.5533546854802523.553\,354\,685\,480\,2523.553 354 685 480 252
(c) Rightmost tree runtime statistics in nano seconds (ns)
Gibbon Marmoset
lr rl Mgreedy Msolver
Mean 383.2 318.1 312.2 315.4
Median 360.0 300.0 300.0 300.0
ub 404.5 328.9 321.1 327.1
lb 362.4 307.2 303.4 303.8
5.2.2 Short-circuiting logical And/Or operators
These microbenchmarks operate over a synthetically generated, balanced syntax-tree with a height of 30. The intermediate nodes can be one of Not, Or, or And, selected at random, and the leaves consist of booleans wrapped in a Val. The evaluator function evaluates the expression tree and performs short circuiting for expressions like and and or in the expression tree. We measure the performance of this evaluator function by changing the order of the left and right subtrees. The syntax-tree datatype is defined as follows:
@@ 5.2.2 Short-circuiting logical And/Or operators
Since the short circuiting evaluates from left to right order of the Exp, changing the order of the left and right subtrees would affect the performance of the traversal. As can be seen in table 1(b), the layout where the left subtree is serialized before the right subtree results in better performance compared to the tree where the the right subtree is serialized before the left one. This is as expected since in the latter case, the traversal has to jump over the right subtree serialized before the left one in order to evaluate it first and then depending on the result of the left subtree possibly jump back to evaluate the right subtree. This results in poor spatial locality and hence worse performance. Mgreedy and Msolver are able to identify the layout transformations that would give the best performance, which matches the case where the left subtree is serialized before the right subtree. Table 1(b) [^8] The layout chosen by Mgreedy and Msolver is same as lr, the performance differs from the lr layout compiled through Gibbon as Marmoset does code motion which results in different code. shows the performance results.

### v3

@@ Abstract
While programmers know that memory representation of data structures can have significant effects on performance, compiler support to optimize the layout of those structures is an under-explored field. Prior work has optimized the layout of individual, non-recursive structures without considering how collections of those objects in linked or recursive data structures are laid out.
@@ 2.2 Running Example
A non-empty blog value stores a content field (a string representing the body of the blog post), a list of hash tags summarizing keywords of the blog post, and a pointer to the rest of the list. [^2] In practice, you may want to reuse a standard list type, e.g.: type BlogList = [Blog] data Blog = Blog Content HashTags but a typical compiler (including Gibbon) would specialize the parametric list type with the Blog type to arrive at an equivalent of the definition shown in the main text above.. The datatype has one point of recursion and several variable-length fields in the definition. To extend on this, Section 5.3 contains an example of tree-shaped data (two points of recursion, in particular) with a fixed-length field. The most general case of multiple points of recursion and variable-length fields is also handled by Marmoset.
@@
Marmoset, described in the following section, automates finding the weaknesses of the pointer-based approach and improving data layout and code accordingly. For instance, performance in our example can be improved by swapping the ordering between the Content and HashTags fields. Given this reordering, the hash tags are available directly at the start of the value, which lines up with the algorithm better, as the algorithm always accesses this field first. Additionally, our program’s True case (the keyword gets a hit within the hash tags) is more efficient because after traversing the content to highlight the keyword it stops at the next blog entry ready for the algorithm to make the recursive call. This improved data layout results in the more-streamlined access patterns shown in Figures 3(b) and 3(d).
@@ 3.1 Marmoset’s Language
Marmoset operates on the language $\lambda_{M}$ given in Figure 4.
$\lambda_{M}$ is a first-order, monomorphic, call-by-value functional language with algebraic datatypes and pattern matching. Programs consist of a series of datatype definitions, function definitions, and a main expression. $\lambda_{M}$’s expressions use A-normal form [12]. The notation $\overline{x}$ denotes a vector $[x_{1},\ldots,x_{n}]$ and $\overline{x_{\mathit{i}}}$ the item at position $\mathit{i}$. $\lambda_{M}$ is an intermediate representation (IR) used towards the front end in the Gibbon compiler. The monomorphizer and specializer lower a program written in a polymorphic, higher-order subset of Haskell [^3] With strict evaluation semantics using -XStrict. to $\lambda_{M}$, and then location inference is used to convert it to the location calculus (LoCal) code next [23]. It is easier to update the layout of types in $\lambda_{M}$ compared to LoCal, as in $\lambda_{M}$ the layout is implicitly determined by the ordering of fields, whereas the later LoCal IR makes the layout explicit using locations and regions (essentially, buffers and pointer arithmetic).
@@ 3.2 Control-Flow Analysis
We construct a control-flow graph with sub-expressions, and let-bound RHS’s (right hand sides) of $\lambda_{M}$ as the nodes. Algorithm 1 shows the psuedocode for generating the control-flow graph. Because the syntax is flattened into A-normal form, there is no need to traverse within the RHS of a let expression. Edges between the nodes represent paths between expressions. The edges consist of weights (Line 11) that represent the likelihood of a particular path being taken. An edge between two nodes indicates the order of the evaluation of the program. A node corresponding to a let-binding (Line 9) contains the bound expression and has one outgoing edge to a node corresponding to the body expression. A case expression (Line 14) splits the control flow $n$-ways, where $n$ is the number of pattern matches. Outgoing edges of a node for a case expression have weights associated with them that correspond to the likelihood of taking a particular branch in the program. Control flow terminates on a leaf $\lambda_{M}$ expression: a variable reference, a data constructor or a function application.
Figure 5(a) shows the control-flow graph for the running example (Figure 1). Each node corresponds to a sub-expression of the function emphKeyword. The first case expression splits the control flow into two branches, corresponding to whether the input list of blogs is empty or not. The branch corresponding to the empty input list is assigned a probability $\alpha$, and the other branch is assigned a probability $1-\alpha$. The next node corresponds to the pattern match Blog content hashTags blogs'. Another two-way branch follows, corresponding to whether keyword occurs in the content of this blog or not. We assign the probabilities $\sigma$ and $1-\sigma$ to these branches respectively. Note that as a result, the corresponding edges in the CFG have weights $(1-\alpha)*\sigma$ and $(1-\alpha)*(1-\sigma)$, as the likelihood of reaching that condition is $(1-\alpha)$. Each of these branches terminate by creating a new blog entry with its content potentially updated. In the current model, $\alpha$ and $\sigma$ are 0.5: they are uniformly distributed.
@@
Figure 5: Control-flow and corresponding field access graphs generated for the running example.
@@ 3.4 Field Attributes For Code Motion
When trying to find the best layout, we may treat the code as immutable, but allowing ourselves to move the code around (i.e. change the order of accesses to the fields) unlocks more possibilities for optimizing layout. Not all code motions are valid due to data dependencies in the traversal. For instance, in a sequence of two let binders, the second one may reference the binding introduced in the first one: in this case, the two binders cannot be reordered.
@@
A single field can have multiple attributes. For instance, a function traverse doing a pre-order traversal on a Tree makes recursive calls on both the left and right children. The left and right children are recursive and self-recursive. Therefore, when looking at the scope of traverse, the left and right children have the attributes recursive, self-recursive and inlineable.
Table 1: Running times for two traversals working with their corresponding matching layouts.
datatype / traversal List / foo List' / foo'
1M elements 0.992s 0.953s
@@
Example. Consider the example of a list traversal shown in Figure 6(a). Here the function foo does some work on the Int field (it raises the Int to the power of 100) and then recurs on the tail of the list. Since Marmoset compiles to dense representations, a List’s representation in memory stores the Cons tag (one byte) followed by the Int (8 bytes) followed by the next Cons tag and so on. Hence, the Cons tag and the Int field are interleaved together in memory. The function foo becomes a stream processor that consumes one stream in memory and produces a dense output buffer of the same type.
Alternatively, another layout of a list follows from the following definition:
@@
In fact, we made a similar program and ran Marmoset on the program in the hope that it would be able to optimize the program by placing the inlineable field first and use code-motion to change the traversal. We show the results in table 1. We see a speedup of 4.1% over our baseline datatype and traversal.
@@ 3.5 Field Access Pattern Analysis
18:             mutate edges'= addEdge(edges', ((lastAccessedVar, var), weight), ControlFlowTag)
@@
20:                 mutate edges'= addEdge(edges', ((lastAccessedVar, var), weight), DataFlowTag)
@@ 3.6.4 Greedy layout ordering
Finding an optimal layout using an external solver hurts compile times. To solve this tradeoff, we propose a simple algorithm that traverses the field access graph in a greedy fashion. The algorithm starts from the root node of the graph, which corresponds to the field accessed first in the function, and greedily visits the child nodes based on the edge weights. We fix the edge order for a control-flow edge as the original order and do not look at field attributes. However, after the greedy algorithm picks a layout we match the let expressions to the layout order to make sure the code matches the layout order. The greedy algorithm is potentially sub-optimal when it comes to finding the best performing layout; however, the compile time is fast.
@@ 3.8 Finding a layout for functions with conflicting access patterns
If two functions, f1 and f2, access the fields of D in the opposite orders, we get conflicting access patterns for D. For instance, assume f1 accesses A first and then B, while f2 accesses B first and then A. After combining edges across the two functions, we get two edges in opposition to each other. Since we use uniform weights for all functions, the edges will also have a uniform weight. As a result, placing A before B or vice versa are equally good in our cost model, and Marmoset defers to the solver to get one of the two layouts.
With the two equally good layouts, Marmoset’s solver (the default mode) chooses the layout favoring the function it picked first. For instance, if f1 is defined earlier in the program, the order favoring f1 will be picked. In the greedy mode, since both A and B are root nodes, Marmoset will pick the first root node in the list of root nodes, which is, again, dependent on the ordering of functions in the source code.
@@ 5 Evaluation
We evaluate Marmoset on three applications. First is a pair of microbenchmarks (Section 5.2) —a list length function and a logical expression evaluator— that help us explore performance penalties imposed by a sub-optimal data layout. Second is a small library of operations with binary trees (Section 5.3). Third is a blog management software based on the BlogList example from the Sections 23 (Section 5.4). Besides the run times, we take a closer look at how Marmoset affects cache behavior (Section 5.5) and compile times (Section 5.6). Finally, we discuss evaluation and its scale (Section 5.7).
We detail the impact of various datatype layouts on the performance. As the baseline, we use Gibbon, the most closely related prior work. We also compare Marmoset with MLton (Section 5.4.2). For each benchmark, we run 99 iterations and report the run-time mean and the 95% confidence interval.
@@ 5.2 Micro Benchmarks
ListLength
This benchmark computes the length of a linked-list and demonstrates the cost of de-referencing memory addresses that are not present in the cache. It uses the linked list datatype:
@@
In our experiment, the linked list is made of 3M elements and each element contains an instance of the Pandoc Inline datatype that occupies roughly 5KB. As seen in table 2777The performance of List' layout compiled with Gibbon differs from Marmoset as code motion to reorder let expressions results in different code. In addition to a noisy server., the performance of the list constructed using the original List is $\sim$42$\times$ worse than the performance with the Marmoset-optimized, flipped layout. Not only does List have poor data locality and cache behavior, but it also has to execute more instructions to de-reference the pointer. Both Mgreedy and Msolver choose the flipped layout List'.
Table 2: Run-time mean and 95% confidence interval (ub, lb) for different layouts (seconds). The last two columns show the run time for the layouts chosen by Mgreedy and Msolver. The numbers in blue correspond to the lowest running time and the numbers in red correspond to the highest running time. Legend: l – left subtree, r – right subtree of the tree.
Benchmark name Gibbon Marmoset
@@
ListLength 62.34 (62.26, 62.41) 1.51 (1.44, 1.59) 1.49 (1.41, 1.56) 1.50 (1.42, 1.58)
@@
LogicEval 4.451 995 070 707 074.451995070707074.451\,995\,070\,707\,074.451 995 070 707 07 (4.421 091 032 316 9514.4210910323169514.421\,091\,032\,316\,9514.421 091 032 316 951, 4.482 899 109 097 1884.4828991090971884.482\,899\,109\,097\,1884.482 899 109 097 188) 6.597 977 737 373 7386.5979777373737386.597\,977\,737\,373\,7386.597 977 737 373 738 (6.587 444 262 502 3356.5874442625023356.587\,444\,262\,502\,3356.587 444 262 502 335, 6.608 511 212 245 141 56.60851121224514156.608\,511\,212\,245\,141\,56.608 511 212 245 141 5) 3.555 498 060 606 059 83.55549806060605983.555\,498\,060\,606\,059\,83.555 498 060 606 059 8 (3.533 075 232 550 711 83.53307523255071183.533\,075\,232\,550\,711\,83.533 075 232 550 711 8, 3.577 920 888 661 4083.5779208886614083.577\,920\,888\,661\,4083.577 920 888 661 408) 3.551 578 959 595 959 73.55157895959595973.551\,578\,959\,595\,959\,73.551 578 959 595 959 7 (3.550 932 274 913 1443.5509322749131443.550\,932\,274\,913\,1443.550 932 274 913 144, 3.552 225 644 278 775 63.55222564427877563.552\,225\,644\,278\,775\,63.552 225 644 278 775 6)
Rightmost 384.4 (368.4, 400.3) 314.5 (303.7, 325.3) 306.9 (295.6, 318.2) 303.1 (292.6, 313.6)
LogicEval
This microbenchmark implements a short-circuiting logical expressions evaluator and runts it over synthetically generated, balanced syntax-trees with the height of 30. The intermediate nodes can be one of Not, Or, or And, selected at random, and the leaves hold boolean values. The syntax-tree datatype is defined as follows:
@@
We measure the performance of the evaluator for differeent orders of the left and right subtrees. Since the short circuiting evaluates from left to right order of the Exp, changing the order of the left and right subtrees would affect the performance of the traversal. As can be seen in Table 2, the layout where the left subtree is serialized before the right subtree results in better performance compared to the tree where the right subtree is serialized before the left one. This is as expected since in the latter case, the traversal has to jump over the right subtree serialized before the left one in order to evaluate it first and then depending on the result of the left subtree possibly jump back to evaluate the right subtree. This results in poor spatial locality and hence worse performance. Mgreedy and Msolver are able to identify the layout transformations that would give the best performance, which matches the case where the left subtree is serialized before the right subtree (Table 2888The layout chosen by Mgreedy and Msolver is same as lr, the performance differs from the lr layout compiled with Gibbon as Marmoset does code motion which results in different code.).